A pendulam has time period 4/5 if we give force f1 to it and time period 3/5 if we give force f2 to it .then find time period when we give simultaneously both f1f2 to it in same direction

Grade:12

1 Answers

Vikas TU
4 years ago
Dear student 
Force F = mrw^2 
= mr(2*pi/T)^2 =4*pi^2 mr/T^2 
As te forces re acting simutaneously 
F = F1 + F2 
1/T^2 = 1/T1^2 + 1/T2^2 
Because m and r are constant 
Put the value of T1 and T2 
1/T = 12/25 sec 
Hope this helps 
Good Luck